The letters of the word Mississippi are placed in a hat. What is the probability that the first two letters drawn are both i if the first letter drawn is not replaced?
Answer by nerdybill(7384)

The letters of the word Mississippi are placed in a hat. What is the probability that the first two letters drawn are both i if the first letter drawn is not replaced?
.
first pick:
Total letters: 11
i's : 4
probability 4/11
.
second pick:
Total letters: 10
i'2 : 3
probability 3/10
.
4/11 * 3/10 = 12/110 = 0.11 or 11%